A common flashlight bulb is rated at 0.30A0.30A and 2.9V2.9V (the values of the current and voltage under operating conditions). If the resistance of the tungsten bulb filament at room temperature (20°C)(20°C) is 1.3 ΩΩ, what is the temperature in degrees Celsius of the filament when the bulb is on?
Answer #1

Given that when bulb operate,

Current through bulb = 0.30 A

V = Voltage = 2.9 V

So, Using ohm's law:

V = I*R

R = V/I = 2.9/0.30

R = 9.67 Ohm

Now also given that, at 20 C, resistance R0 = 1.3 Ohm

We know that relation between resistance and temperature is given by:

R = R0*(1 + alpha*(T - T0))

alpha*R0*(T - T0) = R - R0

T = T0 + (R - R0)/(alpha*R0)

Using known values:

alpha = temperature coefficient for tungsten = 4.5*10^-3  

T = 20 + (9.67 - 1.3)/(4.5*10^-3*1.3)

T = 1450.77 C = Temperature of filament when bulb is on

Please Upvote.